7. Suppose that at a price of $70 the quantity supplied in a market is 10 units, and at a price of s80 th e quantity supplied in the market is 15 unit. If we use this information to create a linear supply equation, what will that equation be? b. P-50+ 2Qs Suppose that college tuition is higher this year than last year and that more students are enrolled in college this year than last year. Based on this information, we can best conclude that 8. The law of demand is invalid Despite the increase in price, enrollment increased due to some other factor(s) changing in the market The demand for a college education is positively sloped Everything you have ever believed in has been a lie a. b. c. d. Price Quantity Demanded Quantity Supplied $5 $4 $3 $2 $1 Per Month 500 700 850 1000 1200 Per Month 1200 1000 850 700 500 Consider the table above.If the price in the market is initially set at $2, what is the result in the market, and what will eventually have to happen to move the market to equilibrium? 9. a. Shortage, price increase b. Shortage, price decrease c. Surplus, price increase d. Surplus, price decrease 10. Suppose a market is initially in equilibrium. Then a change occurs and the equilibrium price What change occurred in the market to decreases while the equilibrium quantity increases. cause these changes to price and quantity? a. b. c. d. Decrease in demand Increase in demand Decrease in supply Increase in supply 11. Suppose the market for a good is initially in equilibrium, and then a shock occurs which causes the demand to increase. Which of the following responses do we expect in the market? a. b. c. d. A movement to the left along the supply curve A movement to the right along the supply curve The supply curve will shift left The supply curve will shift right
Answer #1

7. B) P= 50 +2Qs

Case 1: Q=10

P= 50+2*10= 50+20=70

Case 2: Q=15

P= 50+2*15= 50+30=80

8.c. the demand for a college education is positively slopes.

9.shortage; price increase.

at $2, demand =1000, supply=700.

As demand is more than supply, there will be be a shortage and price should go up to maintain equilibrium.

10.c. decrease in supply.

11.the supply curve will shift right

In order to fulfill the demand.